### v1

@@ 2.3 Neutral atomic hydrogen (Hi) emission
We used the data distributed in FITS-format binary tables containing lists of spectra sampled on a HEALPix grid with $N_{\rm side}$ = $1024$, which corresponds to a pixel angular size of 3.44 (Górski et al., 2005). We arranged these spectra in all-sky HEALPix maps corresponding each of the 1.29-km s-1-wide velocity channels. Using the tools in the Python healpy package, we produced a Cartesian projection for each one of these velocity channels into the same spatial grid used for the 3D dust data. An example of the resulting Hi emission data is shown on the bottom panel of Fig. 1.
Figure 2: Rendering of the 3D dust density distribution from the extinction models presented in (Edenhofer et al., 2024) for the region $|b|$$<$ 5 considered in this paper. The yellow sphere represents the position of the Sun. For an online interactive version of the figure, visit https://solerjuan.github.io/astrosoler/visual/soler_2024_interactive.html.
@@ 3.1 Gradient comparison
We matched the Hi and CO emission across velocity channels and the 3D dust cube using the histogram of oriented gradients (HOG) method introduced by Soler (2019). This method is based on characterizing the similarities in the emission distribution using the orientation of its gradients. In the HOG method, a LOS-velocity channel map from the line emission and a distance channel from the 3D dust are morphologically similar if their gradients are mainly parallel and dissimilar if they are randomly oriented. The distribution of angles between gradient vectors is evaluated using the projected Rayleigh statistic ($V$), a statistical test of non-uniformity in an angle distribution around a particular direction (see Jow et al., 2018, and references therein).
We implemented HOG as follows. We split the sky band within $|b|$$<$ 5 into 36 10$\times$ 10 regions, identified by the index $k$. For each region, we used the PPV cube derived from the line emission by tracer $X$, $I_{ijkp}^{\rm X}$, and the nucleon density PPD cube derived from the 3D dust reconstruction $n_{ijkq}$. The $i$ and $j$ indexes run over the pixels in the sky coordinates, which in our case are Galactic longitude ($l$) and latitude ($b$) and the indexes $p$ and $q$ run over LOS velocity and distance channels, respectively. We calculated the relative orientation angles between the emission and density gradients,
$\theta^{X}_{ijkpq}=\arctan\left(\frac{(\nabla I^{\rm X}_{ijkp}\times\nabla n_{ijkq})\cdot\hat{z}}{\nabla I^{\rm X}_{ijkp}\cdot\nabla n_{ijkq}}\right),$ (1)
where $\nabla$ is the differential operator corresponding to the gradient in the spatial sky coordinates $l$ and $b$.
@@ 3.2 Quantifying the morphological correlation
Figure 4: Example of the HOG method morphological correlation between the 3D dust density reconstruction and the line emission observations for the region 30$<$$l$$<$ 40 and $|b|$$<$ 5. Each panel corresponds to the correlation between distance channels, from the 3D dust density, and velocity channels, from the Hi and CO line emission, shown in the left and right panels, respectively. The correlation metric is the direction-sensitive projected Rayleigh statistic ($V_{\rm d}$, Eq. 2). Values of $V_{\rm d}$$\approx$ 0 correspond to a random orientation between the gradients of the two tracers, thus indicating low morphological correlation. Values of $V_{\rm d}$$>$ 2.87 correspond to mostly parallel gradients, thus indicating a significant morphological correlation. Values of $V_{\rm d}$$<$ 2.87 correspond to mostly antiparallel gradients. The white markers indicate the $v_{\rm LOS}$ with the highest $V_{\rm d}$ for each distance channel. The green dashed lines correspond to the distance and $v_{\rm LOS}$ with the highest $V_{\rm d}$ for the whole velocity and distance ranges. The three cyan lines represent the kinematic distances for the Galactic longitude at the center and at the region’s edges, according to the Reid et al. (2019) Galactic rotation model.
@@ 3.2.1 Selection of HOG parameters
3.2.2 Error estimation
We estimated the errors in calculating $V_{\rm d}$ by applying the HOG method to the line emission cubes and the twelve 3D dust distribution realizations produced by Edenhofer et al. (2024) from the variational posterior. The result is twelve $V_{\rm d}$ values for each line emission and 3D dust channel map. Additionally, we analyzed 100 realizations of the line emission maps produced with Monte Carlo sampling, assuming a Gaussian distribution centered on the observed value in each channel and a standard deviation equal to the noise level of each data set. We report the mean value of $V_{\rm d}$ for the twelve 3D dust cubes and the 100 realizations, excluding $V_{\rm d}$ values below the standard deviation of this resulting set. Further details on the propagation of the measurement errors and the uncertainties in the 3D dust reconstruction into the HOG results are discussed in App. A.2.
@@ 3.3 Dissecting the nearby Galactic plane
We carried out the Galactic plane analysis by splitting the observations within the $|b|$$<$ 5 range into 36 contiguous 10$\times$ 10 tiles centered on $b$$=$ 0. For each tile, we computed Eq. (2) using the HOGcorr_ima routine in the publicly-available astroHOG package [^3] https://github.com/solerjuan/astroHOG. The result is a $V_{\rm d}$ matrix representing the correlation between each distance and velocity-channel map toward a tile, as illustrated in Fig. 4.
@@
We employed the $V_{\rm d}$ matrices to assign a $v_{\rm LOS}$ to each distance channel in a tile. We used the maximum between the critical value 2.87 and the $V_{\rm d}$ standard deviation, $(\sigma_{V_{\rm d}})_{vk}$, as significance threshold for the LOS velocity assignment to each distance channel. If the maximum value of $V_{\rm d}$ is below either of these thresholds, no $v_{\rm LOS}$ is assigned to a distance channel. If more than one velocity channel has $V_{\rm d}$ above these thresholds, we assign the $v_{\rm LOS}$ corresponding to the maximum $V_{\rm d}$. This simplification of the velocity field aims to identify a representative LOS velocity for the bulk of the dust at each distance, and it usually leads to an unambiguous $v_{\rm LOS}$.
Figure 4 illustrates an example of the LOS velocity selection from a $V_{\rm d}$ matrix, where the white markers represent the assigned $v_{\rm LOS}$ for each distance channel. In this particular example, we found that the representative LOS motions for CO roughly follow the expected $v_{\rm LOS}$ from Galactic rotation. In contrast, the Hi $v_{\rm LOS}$ shows deviations up to tens of km s-1 from the CO LOS velocities and the LOS velocities expected from the Galactic rotation model. In the following sections, we discuss this and other dynamical effects on the reconstructed velocity field after presenting the global HOG results.
@@ 4.1 Correlation between 3D dust and line emission morphology
An alternative representation of the correlation between the line emission and the 3D dust is presented in Fig. 6, where we show the maximum values of $V_{\rm d}$ across distances, $\max(V_{\rm d})_{d}$, for the $v_{\rm LOS}$ channels in the 36 tiles in which we divided the $|b|$$<$ 5 sky region. That is, each pixel in Fig. 6 corresponds to the value of Eq. (2) for the particular case of $q$$=$$q^{*}$, where $q^{*}$ is the distance channel with the maximum value of $V_{\rm d}$. Figure 6 can be understood as a different projection of the PPDV space mapped with the HOG method. This projection complements Fig. 5 by showing whether a line emission channel has a morphological match in the 3D dust distribution between 69 and 1250 pc.
@@
The Hi emission in Fig. 7 shows the typical shadows produced by the cold Hi observed on top of the warm background emission, which is a characteristic signature of a HISA feature (see, for example, Gibson et al., 2000). We presume that the lower Hi emission within the CO contours in the central panel of Fig. 7 is a HISA based on the morphological similarity between the CO emission and HISA identified in other observations of the Galactic plane (see, for example, Gibson et al., 2005; Soler et al., 2019; Wang et al., 2020b). However, a confirmation of the HISA for this example and throughout the Galactic plane requires further analysis of the Hi spectra, which is beyond the scope of this paper (see, for example Syed et al., 2023). The presumed HISA in this particular example does not produce a prominent $V_{\rm d}$$<$ 0 in the HOG comparison between Hi and 3D dust, as shown in Fig. 4. This is most likely due to the combination of antiparallel gradients in the HISA feature and the parallel gradients outside it, resulting in $V$$>$ 0.
Figure 8 shows the Hi and CO line emission and density distribution for the tiles with the highest $V_{\rm d}$ in the Hi and 3D dust comparison, as identified in the right-hand-side panel of Fig. 5. In contrast with the region in Fig. 7, there is no CO counterpart to the matching structures in Hi and 3D dust. This example illustrates that the morphological correlation traced by Hi and CO is not necessarily identical because of the different structures each tracer shows.
@@
We further compared the morphological correlations of the 3D dust and the two gas tracers in Fig. 9. The scatter plot contrasts the values of $\max(V_{\rm d})_{v}$ for Hi and CO reported in Fig. 5. We found examples of tiles where the highest $\max(V_{\rm d})$ values are found in the two gas tracers. However, in general, the correlation between the 3D dust and one gas tracer does not indicate its correlation with the other.
We find high $\max(V_{\rm d})_{v}$ for either tracer across multiple distances in Fig. 9. We interpret this observation as an indication that there is no significant bias introduced by distance in the morphological correlation, as further considered in App. C. The highest $\max(V_{\rm d})_{v}$ is found for Hi, as expected from the largest angular extend of Hi emission.
Fig. 9 shows a few tiles where $\max(V_{\rm d})_{v}$ is negative, indicating that despite our selection of the maximum $V_{\rm d}$ in each tile, there are regions where the gradients are mostly antiparallel. In CO, the negative $\max(V_{\rm d})_{v}$ values are a factor of two the significance threshold, $V_{\rm d}$$<$$-2.87$. In HI, there are a few tiles in the $V_{\rm d}$$<$$-2.87$ range, which we further consider in the context of potential HISA detection in App. B.
The low morphological correlation, $V_{\rm d}$$<$ 2.87, can result from several local conditions. First, distance channels with low dust densities would produce gradients dominated by the smoothness priors and uncertainties in the 3D dust extinction reconstruction, which do not correlate with the line emission distribution. These low-density regions are evident in the mean dust density distribution reported in Fig. 3. Second, velocity channels with low line intensities would produce gradients dominated by observation noise, which do not correlate with the structures traced by extinction. This can be due to low gas densities or the prevalence of structures not traced by the gas emission, as discussed in Sec. 6.1. This effect is apparent towards Galactic longitudes in the range $200$$<$$l$$<$ 250 and velocities 0 $<$$v_{\rm LOS}$$<$ 10 km s-1: we see no significant correlation between the dust and CO in Fig. 6 simply because there is little gas or dust present in this region, which is associated with the GSH 238+00+09 superbubble (Heiles, 1998). Finally, low $|V_{\rm d}|$ can result from distance and velocity channels with significant resolved structures that are simply not correlated.
@@ 4.2 Local ISM motions
The streaming motions distribution, shown in Fig. 12, is not strictly symmetric but is centered around zero, indicating that, on average, the departures from pure Galactic rotation are relatively small. The mean values streaming motions are roughly 0.1 and 1.0 km s-1 for Hi and CO, respectively. These values are below the 1.3-km s-1 spectral resolution of the line emission data, so they are within the LOS velocity determination uncertainties. The characteristic amplitude of the streaming motions, quantified by the standard deviation of $v^{\rm R19}_{\rm LOS}-v^{\rm HOG}_{\rm LOS}$, is roughly 12.1 and 8.8 km s-1, for Hi and CO, respectively.
@@
The streaming motions map, presented in Fig. 13, does not show a prevalent large-scale pattern of streaming motions that may indicate the effect of compression toward the Local Arm. Nor does it reveal global agreement in the radial velocity pattern sampled by the Hi and CO line emission, indicating that the residual motions due to Galactic rotation are relatively small. In some positions, for example, at $d$$\sim$ 800 pc for 80 $<$$l$$<$ 90, the Hi shows diverging motions that the CO does not match. However, in general, the radial velocities do not reveal the concentration of converging or diverging motions but rather a succession of compressions and rarefactions along the line of sight, as expected from the behavior of a turbulent medium with an energy cascade encompassing the scales considered in this study.

### v2

@@ 2.3 Neutral atomic hydrogen (Hi) emission
We used the data distributed in FITS-format binary tables containing lists of spectra sampled on a HEALPix grid with $N_{\rm side}$ = $1024$, which corresponds to a pixel angular size of 3.44 (Górski et al., 2005). We arranged these spectra in all-sky HEALPix maps corresponding to the 1.29-km s-1-wide velocity channels. Using the tools in the Python healpy package, we produced a Cartesian projection for each one of these velocity channels into the same spatial grid used for the 3D dust data. An example of the resulting Hi emission data is shown on the bottom panel of Fig. 1.
Figure 2: Rendering of the 3D dust density distribution from the extinction models presented in Edenhofer et al. (2024) for the region $|b|$$<$ 5 considered in this paper. The yellow sphere represents the position of the Sun. The associated movie is available online.
@@ 3.1 Gradient comparison
We matched the Hi and CO emission across velocity channels and the 3D dust cube using the histogram of oriented gradients (HOG) method introduced in Soler et al. (2019). This method is based on characterizing the similarities in the emission distribution using the orientation of its gradients. In the HOG method, a LOS-velocity channel map from the line emission and a distance channel from the 3D dust are morphologically similar if their gradients are mainly parallel and dissimilar if they are randomly oriented. The distribution of angles between gradient vectors is evaluated using the projected Rayleigh statistic ($V$), a statistical test of non-uniformity in an angle distribution around a particular direction (see Jow et al., 2018, and references therein).
We implemented HOG as follows. We split the sky band within $|b|$$<$ 5 into 36 10$\times$ 10 regions, identified by the index $k$. For each region, we used the PPV cube derived from the line emission by tracer $X$, $I_{ijkp}^{\rm X}$, and the nucleon density PPD cube derived from the 3D dust reconstruction $n_{ijkq}$. The $i$ and $j$ indexes run over the pixels in the sky coordinates, which in our case are Galactic longitude ($l$) and latitude ($b$), and the indexes $p$ and $q$ run over LOS velocity and distance channels, respectively. We calculated the relative orientation angles between the emission and density gradients,
$\theta^{X}_{ijkpq}=\arctan\left(\frac{\left\|\nabla I^{\rm X}_{ijkp}\times\nabla n_{ijkq}\right\|}{\nabla I^{\rm X}_{ijkp}\cdot\nabla n_{ijkq}}\right),$ (1)
[…]
@@ 3.2 Quantifying the morphological correlation
Figure 4: Example of the HOG method morphological correlation between the 3D dust density reconstruction and the line emission observations for the region 30$<$$l$$<$ 40 and $|b|$$<$ 5. Each panel corresponds to the correlation between distance channels, from the 3D dust density, and velocity channels, from the Hi and CO line emission, shown in the left and right panels, respectively. The correlation metric is the direction-sensitive projected Rayleigh statistic ($V_{\rm d}$, Eq. 2). Values of $V_{\rm d}$$\approx$ 0 correspond to a random orientation between the gradients of the two tracers, thus indicating low morphological correlation. Values of $V_{\rm d}$$>$ 2.87 correspond to mostly parallel gradients, thus indicating a significant morphological correlation. Values of $V_{\rm d}$$<$ 2.87 correspond to mostly antiparallel gradients. The white markers indicate the $v_{\rm LOS}$ with the highest $V_{\rm d}$ for each distance channel. The green dashed lines correspond to the distance and $v_{\rm LOS}$ with the highest $V_{\rm d}$ for the whole velocity and distance ranges. The three cyan lines represent the kinematic distances for the Galactic longitude at the center and the region’s edges, according to the Reid et al. (2019) Galactic rotation model.
@@ 3.2.1 Selection of HOG parameters
3.2.2 Handling of errors in the HOG morphological correlation and LOS velocity estimation
We estimated the errors in $V_{\rm d}$ by applying the HOG method to the line emission cubes and the twelve 3D dust distribution realizations produced by Edenhofer et al. (2024) from the variational posterior. The result is twelve $V_{\rm d}$ values for each line emission and 3D dust channel map. Additionally, we analyzed 100 realizations of the line emission maps produced with Monte Carlo sampling, assuming a Gaussian distribution centered on the observed value in each channel and a standard deviation equal to the noise level of each data set. We reported the mean value of $V_{\rm d}$ for the twelve 3D dust cubes and the 100 realizations.
The $V_{\rm d}$ standard deviation estimated from the 12 $\times$ 100 samples, $\sigma_{V_{\rm d}}$, is a measurement of variance in the morphological correlation. In App. A.2, we show that $\sigma_{V_{\rm d}}$ is led by the variance among the twelve 3D dust realizations rather than by the noise in the line emission observations. Given that each of the twelve realizations is equally valid, large $\sigma_{V_{\rm d}}$ value reflects an ambiguity in the 3D dust morphology that limits the significance of the morphological matching with a line emission channel. Thus, we exclude distance-$v_{\rm LOS}$ pairs with $|V_{\rm d}/\sigma_{V_{\rm d}}|$$<$ 3.0 from the calculations of the representative $v_{\rm LOS}$ in each distance channel, as described in Sec. 3.3. We also exclude distance-$v_{\rm LOS}$ pairs with $V_{d}$ below the chance correlation thresholds obtained when flipping one of the input maps in the vertical, horizontal, and diagonal directions, as described in App. A.2.3.
@@ 3.3 Dissecting the nearby Galactic plane
We carried out the Galactic plane analysis by splitting the observations within the $|b|$$<$ 5 range into 36 contiguous 10$\times$ 10 tiles centered on $b$$=$ 0. For each tile, we computed Eq. (2) using the HOGcorr_ima routine in the publicly available astroHOG package [^3] https://github.com/solerjuan/astroHOG. The result is a $V_{\rm d}$ matrix representing the correlation between each distance and velocity-channel map toward a tile, as illustrated in Fig. 4.
@@
We employed the $V_{\rm d}$ matrices to assign a $v_{\rm LOS}$ to each distance channel in a tile. We used the critical value 2.87 and the $V_{\rm d}$ standard deviation, $(\sigma_{V_{\rm d}})_{vk}$, as significance thresholds for the LOS velocity assignment to each distance channel. If the maximum value of $V_{\rm d}$ is below either of these values, no $v_{\rm LOS}$ is assigned to a distance channel. If more than one velocity channel has $V_{\rm d}$ above these thresholds, we assign the $v_{\rm LOS}$ corresponding to the maximum $V_{\rm d}$. The selection by maximum $V_{\rm d}$ excludes distance channels potentially dominated by HISA from the $v_{\rm LOS}$ reconstruction. This choice is motivated by analysis in Appendix B, which shows that a selection based on $|V_{\rm d}|$ instead of $V_{\rm d}$ only favors negative values in a small number of the distance channels and that the interpretation of $V_{\rm d}$$<$ 0 exclusively as the product of HISA is not always correct. We reserve the specific study of HISA using HOG for a subsequent publication.
The simplification of the velocity field in our HOG application aims to identify a representative LOS velocity for the bulk of the dust in each distance slice along a 10$\times$ 10 area. Thus, we estimated the average LOS motions of volumes between 12 $\times$ 12 $\times$ 0.4 pc3 and 220 $\times$ 220  $\times$ 7.0 pc3. We acknowledge that this selection extracts just a segment of the turbulent energy cascade, which connects galactic motions to the smallest scales in the ISM. However, our goal is reconstructing the bulk motions of density structures revealed by the 3D dust model rather than describing the full complexity of the ISM velocity field across scales, which has been considered in other works (see, for example, Elmegreen & Scalo, 2004; Hennebelle & Falgarone, 2012, and references therein).
Figure 4 illustrates an example of the LOS velocity selection from a $V_{\rm d}$ matrix, where the markers represent the assigned $v_{\rm LOS}$ for each distance channel. In this example, the representative LOS motions for CO roughly follow the expected $v_{\rm LOS}$ from Galactic rotation, indicated by the solid lines. In contrast, the Hi $v_{\rm LOS}$ shows deviations up to tens of km s-1 from the CO $v_{\rm LOS}$ and the LOS velocities expected from the Galactic rotation model. In the following sections, we discuss this and other dynamical effects on the reconstructed velocity field after presenting the global HOG results.
@@ 4.1 Correlation between 3D dust and line emission morphology
An alternative representation of the correlation between the line emission and the 3D dust is presented in Fig. 6, where we show the maximum values of $V_{\rm d}$ across distances, $\max(V_{\rm d})_{d}$, for the $v_{\rm LOS}$ channels in the 36 tiles in which we divided the $|b|$$<$ 5 sky region. That is, each pixel in Fig. 6 corresponds to the value of Eq. (2) for the particular case of $q$$=$$q^{*}$, where $q^{*}$ is the distance channel with the maximum value of $V_{\rm d}$. Figure 6 can be understood as a different projection of the PPDV space mapped with the HOG method, although some of the line emission signal in the $|v_{\rm LOS}|$$\leq$ 25 km s-1 may come from density structures beyond the 3D dust reconstruction distance range. Therefore, Fig. 6 should be understood as a visual representation of whether a line emission channel has a morphological match in the 3D dust distribution between 69 and 1250 pc, rather than as a full reconstruction of the PPV space from PPD information.
@@
The Hi emission in Fig. 7 shows the typical shadows produced by the cold Hi observed on top of the warm background emission, which is a characteristic signature of a HISA feature (see, for example, Gibson et al., 2000). We presume that the lower Hi emission within the CO contours in the central panel of Fig. 7 is a HISA based on the morphological similarity between the CO emission and HISA identified in other observations of the Galactic plane (see, for example, Gibson et al., 2005; Soler et al., 2019; Wang et al., 2020). However, a confirmation of the HISA for this example and throughout the Galactic plane requires further analysis of the Hi spectra, which is beyond the scope of this paper (see, for example Syed et al., 2023). The presumed HISA in this particular example does not produce a prominent $V_{\rm d}$$<$ 0 in the HOG comparison between Hi and 3D dust, as shown in Fig. 4. This is most likely due to the combination of antiparallel gradients in the HISA feature and the parallel gradients outside it, resulting in $V$$>$ 0.
Figure 8 shows the Hi and CO line emission and density distribution for the tiles with the highest $V_{\rm d}$ in the Hi and 3D dust comparison, as identified in the left-hand-side panel of Fig. 5. In contrast with the region in Fig. 7, there is no CO counterpart to the matching structures in Hi and 3D dust. This example illustrates that the morphological correlation traced by Hi and CO is not necessarily identical because of the different structures each tracer shows.
@@
We further compared the morphological correlations of the 3D dust and the two gas tracers in Fig. 9. The scatter plot contrasts the values of $\max(V_{\rm d})_{v}$ for Hi and CO reported in Fig. 5. There are some tiles where $\max(V_{\rm d})$ is high for the two gas tracers. However, the correlation between the 3D dust and one gas tracer does not indicate its correlation with the other.
We found that the highest $\max(V_{\rm d})_{v}$ are not concentrated around near or far distances. We interpret this observation as an indication that no significant bias is introduced by distance in the morphological correlation, as further considered in App. C. The highest $\max(V_{\rm d})_{v}$ is found for Hi, as expected from the largest angular extend of Hi emission.
Figure 9 shows a few tiles where $\max(V_{\rm d})_{v}$ is negative, most of them in Hi. This suggests that despite our selection of the maximum $V_{\rm d}$ in each tile, there are regions where the Hi emission gradients are mostly antiparallel to the 3D dust gradients. However, their significance is low, $V_{\rm d}$$<$$-2.87$, and they are related to chance correlation rather than HISA, as discussed in App. B.
@@ 4.2 Local ISM motions
Appendix D presents the HOG analysis of synthetic line emission from a simulated MC, demonstrating that our approach correctly reproduces the mean $v_{\rm LOS}$ of the numerical cloud. Dispersions around the central $v_{\rm LOS}$ are dominated by the emission linewidth. Larger velocity excursions registered with the HOG are produced by the MC’s complex dynamic structure rather than by a systematic effect introduced by the method.
@@
The streaming motions distribution, shown in Fig. 12, is not strictly symmetric but is centered around zero, indicating that, on average, the departures from pure Galactic rotation are relatively small. The mean values streaming motions are roughly 0.2 and 1.1 km s-1 for Hi and CO, respectively. These values are below the 1.3-km s-1 spectral resolution of the line emission data, so they are within the LOS velocity determination uncertainties. The characteristic amplitude of the streaming motions, quantified by the standard deviation of $v^{\rm R19}_{\rm LOS}-v^{\rm HOG}_{\rm LOS}$, is roughly 10.8 and 6.6 km s-1, for Hi and CO, respectively. Given the $|v_{\rm LOS}|$$<$ 25 km s-1 input range limitation, large excursions from Galactic rotation are currently excluded from the reconstruction, so these dispersions may represent lower limits of the actual streaming motions.
@@
The streaming motions map, presented in Fig. 13, does not show a prevalent large-scale pattern of streaming motions that may indicate the effect of compression toward the Local Arm. Nor does it reveal global agreement in the radial velocity pattern sampled by the Hi and CO line emission, indicating that the residual motions due to Galactic rotation are relatively small. In some positions, for example, at $d$$\sim$ 800 pc for 80 $<$$l$$<$ 90, the Hi shows diverging motions that the CO does not match. However, in general, the radial velocities do not reveal the concentration of converging or diverging motions but rather a succession of compressions and rarefactions along the line of sight, as expected from the behavior of a turbulent medium. Given the scales considered in this study, averaging of cells that range between 12 pc $\times$ 12 pc $\times$ 0.4 pc and 218 pc $\times$ 218 pc $\times$ 7 pc in size, it is plausible that the reported motions result from a turbulent energy cascade introduced at kiloparsec scales by global galactic motions (see, for example, Colman et al., 2022, and references therein).